OpenStudy (anonymous):

Show me the steps please? Given f(x)=x^2+6x+8 and g(x)=2x-5... Find: (f+g)(-2). Thanks.

OpenStudy (anonymous):

These are functions :)

OpenStudy (ajprincess):

to find (f+g)(x) Simply add f(x) and g(x). That is (f+g)(x)=x^2+6x+8+2x-5 To find (f+g)(-2) Put -2 in place of x in (f+g)(x)

OpenStudy (ajprincess):

Can u do @starshine7?

OpenStudy (anonymous):

I'll try, be right back :)

OpenStudy (anonymous):

i got -17

OpenStudy (ajprincess):

(f+g)(x)=x^2+6x+8+2x-5 =x^2+8x+3 (f+g)(-2)=(-2)^2+8(-2)+3 =4-16+3 =7-16 =9

OpenStudy (ajprincess):

it's -9